OpenStudy (anonymous):

How do I solve n/-1.5 = -1.2?

OpenStudy (phi):

what happens when you multiply both sides by -1.5 ?

OpenStudy (anonymous):

n/-1.5 * -1.2/-1.5 = -1.8/2.25 ..?

OpenStudy (phi):

like this \[ \frac{n}{-1.5} \cdot -1.5 = -1.2 \cdot -1.5 \]

OpenStudy (anonymous):

Ah, okay, what do I do next?